OpenStudy (anonymous):

The Parent Club wants to sell buttons to raise money for the local community center sports program. The buttons cost 60 cents each, and the club buys 1,200 of them. How much total profit will they make if they charge $4 for each button and sell all of the buttons?

OpenStudy (anonymous):



OpenStudy (lynfran):

1200*4=?

OpenStudy (anonymous):

4,800

OpenStudy (anonymous):

do i divide that by 60?

OpenStudy (anonymous):

@LynFran

OpenStudy (lynfran):

no .60*1200=?

OpenStudy (anonymous):

72,000

OpenStudy (lynfran):

no 0.60*1200=?

OpenStudy (anonymous):

720

OpenStudy (lynfran):

so to find the we take the 720 from 4800 ...we get?

OpenStudy (anonymous):

4,080

OpenStudy (lynfran):

correct and that would be the profit they make

OpenStudy (anonymous):

thank you..heres a medal

OpenStudy (lynfran):

thanks

OpenStudy (anonymous):

yw